# LIGHT: Multi-Modal Text Linking on Historical Maps

### v1

@@ 1 Introduction
Linking text (words) in reading order on historical maps is challenging due to their complex and diverse layouts. Figure 1 shows three example maps with annotations where a highlighted color indicates a multi-word location phrase. These annotations present several challenges for automatic linking methods. For example, a place name might not fall on a single straight line, or unrelated text labels may appear between multi-word phrases, e.g., “SAINT HELENA” in Figure 1(c). High-density text regions further complicate distinguishing linked groups, as in Figure 1(a). Additionally, inconsistent layouts across maps can have words far apart representing large regions, e.g., “DENVER AND RIO GRANDE” in Figure 1(b). These challenges require text linking methods to effectively capture word relationships from multiple perspectives, including linguistic, visual, and geometric contexts.
@@
Figure 2: LIGHT preview: Language, Image, and Geometry for Linking Hierarchical Text on maps.
To address these challenges, we propose a novel multi-modal approach called LIGHT, which integrates Linguistic, Image/visual, and Geometric contexts for linking Hierarchical Text on maps. As shown in Figure 2, LIGHT first employs a polygon encoder to transform the polygonal coordinates of text regions into fixed-size vector embeddings, capturing irregular polygon shapes and relative positions of text instances. Then LIGHT utilizes a multi-modal transformer encoder that integrates the transcribed text contents, map image, and polygon embeddings to jointly learn a cross-modal representation for each word. Finally, LIGHT’s successor predictor directly calculates a one-versus-all probability of each word being the successor for a given word, followed by a greedy selection strategy to determine the optimal link sequence. Additionally, we introduce a bi-directional learning strategy that predicts both the successor and predecessor of each word, enhancing the model’s robustness in learning text sequences. The main contributions of the paper include,
@@ 2.2 Document Layout Analysis
3 LIGHT: Language, Image, and Geometry for Linking Hierarchical Text
Figure 3 shows the overall architecture of LIGHT for linking text in reading order on historical maps. Given a map image with the detected and recognized text instances, LIGHT’s polygon encoder first transforms the polygonal coordinates of each text instance into a fixed-size embedding. Then LIGHT integrates the polygon embeddings with the linguistic token embeddings and image features as input to a multi-modal transformer encoder to learn a cross-modal representation for each word. To adapt to historical map data, we pretrain both encoders using text labels and map images. LIGHT leverages the learned representations to predict the successor of each word and construct text links.
Figure 3: LIGHT contains three components: a Polygon Encoder, a Multi-Modal Transformer Encoder, and a Successor Predictor. Dashed purple arrows indicate the pretraining flow. Underlined purple labels are pretraining objectives for the two encoders. Solid pink arrows show the finetuning and inference flows.
@@ 3.1 Polygon Encoder
3.2 Multi-Modal Transformer Encoder
LIGHT employs a multi-modal transformer encoder to generate cross-modal representations for text instances. Existing document analysis methods, such as LayoutLMv3 [8], jointly model text and image but overlook text shapes (or only use bounding boxes to indicate text positions), which is crucial for handling irregularly shaped text on historical maps. LIGHT builds on LayoutLMv3 by incorporating polygon embeddings, i.e., [CLS] token embeddings, generated by the polygon encoder to enrich geometric information.
@@ 3.2 Multi-Modal Transformer Encoder
We pretrain the multi-modal transformer encoder with the polygon encoder on map text and images. To obtain text annotations for pretraining, instead of using the OCR tool in LayoutLMv3, we use a text spotting model specialized in historical maps [20] to detect and recognize text instances on map images, which generates a large amount of training data (details in Section 4.1.2). We follow the pretraining steps outlined in LayoutLMv3, utilizing three objectives: masked language modeling (MLM), masked image modeling (MIM), and word-patch alignment (WPA). For MLM, we randomly mask 30% of the text tokens, and the pretraining objective is to reconstruct the masked tokens based on the contextual representations of the corrupted sequence of text and image tokens and geometric information added to the text tokens. For MIM, we randomly mask 40% of the image tokens, and similarly, the objective is to reconstruct the masked image tokens [2]. We use a pretrained image tokenizer to generate the labels of image tokens, which can transform dense image pixels into a discrete token with a visual vocabulary [36]. For WPA, the intuition is to perform explicit alignment learning between text and image modalities, as the masking processes are independent for MLM and MIM. The objective is to predict whether the corresponding image patches of a text word are masked. By training with these three objectives on map data, the model can capture cross-modal relationships between map text (contents and shapes) and images. LIGHT uses the final token embeddings for the text linking task, denoted as $E_{\text{text}}\in\mathbb{R}^{N\times D}$, where $N$ is the number of words and $D$ is the embedding dimension ($D=768$).
@@ 3.3.3 Link Inference
During inference, LIGHT constructs text linking paths by iteratively selecting the most probable successor for each word in a greedy manner. Given a target word, LIGHT identifies the next word with the highest predicted probability as its successor. However, this process can lead to ambiguities where a word is assigned to multiple predecessors. To ensure a word belongs to only one phrase, LIGHT enforces a selection rule: if a word has multiple predecessors, only the word with the highest probability is retained as the unique predecessor, while other predecessors continue searching for their next-best successor. This greedy strategy ensures that each word is linked to at most one predecessor, preventing conflicts and preserving a clear, coherent reading order for text. Then LIGHT generates sequences from these word-to-successor links. We provide the pseudo-code in the supplementary material.
@@ 4.1.1 Text Linking Dataset
We use the Rumsey benchmark dataset [21, 24] from the ICDAR 2024/2025 MapText Competitions [17, 23] for model training (200 tiles), validation (40 tiles), and testing (700 tiles). Each map tile has a resolution of $000\times 2{,}$ 000 × 2 , 000 pixels.
@@ 4.1.2 Pretraining Dataset
We leverage additional map images from the David Rumsey Map Collection [4] to pretrain LIGHT’s polygon encoder and multi-modal transformer encoder using self-supervised learning. We generate 397,385 cropped map tiles from 41,279 maps with a mix of both $000\times 1{,}$ 000 × 1 , 000 and $000\times 2{,}$ 000 × 2 , 000 pixels. To produce word-level detections and transcriptions, we run a state-of-the-art map text spotting model [20], which was pretrained on synthetic map image and scene image datasets, and finetuned on the MapText Rumsey training data. These automatically generated polygons and text transcriptions serve as labels for pretraining.
@@ 4.2.2 Training Settings
For the polygon encoder, we set the maximum sequence length as 32, the number of layers as 6, and the embedding dimension as 768. Pretraining is conducted for 100,000 steps with a batch size of 8, using a maximum learning rate (LR) of 0.0001, a 10% warm-up phase, and a linear LR decay [12]. The multi-modal transformer encoder adopts 12 layers and a 768-dimensional embedding size, consistent with LayoutLMv3BASE. Pretraining runs for 375,000 steps with a batch size of 4, using a maximum LR of 0.0001, a 4.8% warm-up phase, and linear decay [8]. During finetuning, we use a batch size of 2 and an initial LR of 0.0005, with a performance-based LR scheduler. If validation performance does not improve for 5 epochs, the LR is reduced by 10%; training is terminated after 9 consecutive non-improving epochs. All the models are trained with one NVIDIA A100 (40GB) GPU. [^1] Code and data: https://github.com/knowledge-computing/multimodal-text-linking
@@ 4.4.5 Visualizations
Figure 4: Successful examples with link recall $\geq$90%.
@@ Appendix 0.A Complexity Analysis
Method #Param (×$\times 10^{6}$) Time (×$\times 10^{-1}$s/image)
LIGHT - Polygon Encoder 70.1 1.6
BERTBASE 111.0 1.7
LayoutLMv3BASE 130.4 1.9
LIGHT 197.1 3.4
@@ Appendix 0.B Dataset Statistics
Words 34,518 5,544 128,457
Valid Words 30,563 4,860 111,821
Label Groups 21,205 3,502 78,582
Groups with $>$1 Words 9,772 1,453 36,655

### v2

@@ Abstract
This version of the contribution has been accepted for publication, after peer review but is not the Version of Record and does not reflect post-acceptance improvements, or any corrections. The Version of Record is available online at: https://doi.org/10.1007/XXXXX. Use of this Accepted Version is subject to the publisher’s Accepted Manuscript terms of use https://www.springernature.com/gp/open-research/policies/accepted-manuscript-terms
To appear in Proceedings of the 19th International Conference on Document Analysis and Recognition (ICDAR 2025). Xu-Cheng Yin, Dimosthenis Karatzas, and Daniel Lopresti (Eds.). Springer Nature LNCS. 2025.
@@ 1 Introduction
Linking text (words) in reading order on historical maps is challenging due to their complex and diverse layouts. Figure 1 shows three example maps with annotations where a highlighted color indicates a multi-word location phrase. These annotations present several challenges for automatic linking methods. For example, a place name might not fall on a single straight line, or unrelated text labels may appear between multi-word phrases, e.g., “SAINT HELENA” in Figure 1(a). High-density text regions further complicate distinguishing linked groups, as in Figure 1(b). Additionally, inconsistent layouts across maps can have words far apart representing large regions, e.g., “DENVER AND RIO GRANDE” in Figure 1(c). These challenges require text linking methods to effectively capture word relationships from multiple perspectives, including linguistic, visual, and geometric contexts.
@@
Figure 2: LIGHT preview: Language, Image, and Geometry for Linking Hierarchical Text on historical maps.
To address these challenges, we propose a novel multi-modal approach called LIGHT, which integrates Linguistic, Image/visual, and Geometric contexts for linking Hierarchical Text on maps. As shown in Figure 2, LIGHT first employs a polygon encoder to transform the polygonal coordinates of text regions into fixed-size vector embeddings, capturing irregular polygon shapes and relative positions of text instances. Then LIGHT utilizes a multi-modal transformer that integrates the transcribed text contents, map image, and polygon embeddings to jointly learn a cross-modal representation for each word. Finally, LIGHT’s successor predictor directly calculates a one-versus-all probability of each word being the successor for a given word, followed by a greedy selection strategy to determine the optimal link sequence. Additionally, we introduce a bi-directional learning strategy that predicts both the successor and predecessor of each word, enhancing the model’s robustness in learning text sequences. The main contributions of the paper include,
@@ 2.2 Document Layout Analysis
3 LIGHT
Figure 3 shows the overall architecture of LIGHT for linking text in reading order on historical maps. Given a map image with the detected and recognized text instances, LIGHT’s polygon encoder first transforms the polygonal coordinates of each text instance into a fixed-size embedding. Then LIGHT integrates the polygon embeddings with the linguistic token embeddings and image features as input to a multi-modal transformer to learn a cross-modal representation for each word. To adapt to historical map data, we pretrain both encoders using text labels and map images. LIGHT leverages the learned representations to predict the successor of each word and construct text links.
Figure 3: LIGHT contains three components: a Polygon Encoder, a Multi-Modal Transformer, and a Successor Predictor. Dashed purple arrows indicate the pretraining flow. Underlined purple labels are pretraining objectives for the two encoders. Solid pink arrows show the finetuning and inference flows.
@@ 3.1 Polygon Encoder
3.2 Multi-Modal Transformer
LIGHT employs a multi-modal transformer to generate cross-modal representations for text instances. Existing document analysis methods, such as LayoutLMv3 [8], jointly model text and image but overlook text shapes (or only use bounding boxes to indicate text positions), which is crucial for handling irregularly shaped text on historical maps. LIGHT builds on LayoutLMv3 by incorporating polygon embeddings, i.e., [CLS] token embeddings, generated by the polygon encoder to enrich geometric information.
@@ 3.2 Multi-Modal Transformer
We pretrain the multi-modal transformer with the polygon encoder on map text and images. To obtain text annotations for pretraining, instead of using the OCR tool in LayoutLMv3, we use a text spotting model specialized in historical maps [20] to detect and recognize text instances on map images, which generates a large amount of training data (details in Section 4.1.2). We follow the pretraining steps outlined in LayoutLMv3, utilizing three objectives: masked language modeling (MLM), masked image modeling (MIM), and word-patch alignment (WPA). For MLM, we randomly mask 30% of the text tokens, and the pretraining objective is to reconstruct the masked tokens based on the contextual representations of the corrupted sequence of text and image tokens and geometric information added to the text tokens. For MIM, we randomly mask 40% of the image tokens, and similarly, the objective is to reconstruct the masked image tokens [2]. We use a pretrained image tokenizer to generate the labels of image tokens, which can transform dense image pixels into a discrete token with a visual vocabulary [36]. For WPA, the intuition is to perform explicit alignment learning between text and image modalities, as the masking processes are independent for MLM and MIM. The objective is to predict whether the corresponding image patches of a text word are masked. By training with these three objectives on map data, the model can capture cross-modal relationships between map text (contents and shapes) and images. LIGHT uses the final token embeddings for the text linking task, denoted as $E_{\text{text}}\in\mathbb{R}^{N\times D}$, where $N$ is the number of words and $D$ is the embedding dimension ($D=768$).
@@ 3.3.3 Link Inference
During inference, LIGHT constructs text linking paths by iteratively selecting the most probable successor for each word in a greedy manner. Given a target word, LIGHT identifies the next word with the highest predicted probability as its successor. However, this process can lead to ambiguities where a word is assigned to multiple predecessors. To ensure a word belongs to only one phrase, LIGHT enforces a selection rule: if a word has multiple predecessors, only the word with the highest probability is retained as the unique predecessor, while other predecessors continue searching for their next-best successor. This greedy strategy ensures that each word is linked to at most one predecessor, preventing conflicts and preserving a clear, coherent reading order for text. Then LIGHT generates sequences from these word-to-successor links. We provide the pseudo-code in the supplementary material Section 0.C.
@@ 4.1.1 Text Linking Dataset
We use the Rumsey benchmark dataset [21, 24] from the ICDAR 2024/2025 MapText Competitions [17, 23] for model training (200 tiles), validation (40 tiles), and testing (700 tiles). Each map tile has a resolution of $000\times 2{,}$ 000 × 2 , 000 pixels. We provide the dataset statistics in the supplementary material Section 0.B.
@@ 4.1.2 Pretraining Dataset
We leverage additional map images from the David Rumsey Map Collection [4] to pretrain LIGHT’s polygon encoder and multi-modal transformer using self-supervised learning. We generate 397,385 cropped map tiles from 41,279 maps with a mix of both $000\times 1{,}$ 000 × 1 , 000 and $000\times 2{,}$ 000 × 2 , 000 pixels. To produce word-level detections and transcriptions, we run a state-of-the-art map text spotting model [20], which was pretrained on synthetic map image and scene image datasets, and finetuned on the MapText Rumsey training data. These automatically generated polygons and text transcriptions serve as labels for pretraining.
@@ 4.2.2 Training Settings
For the polygon encoder, we set the maximum sequence length as 32, the number of layers as 6, and the embedding dimension as 768. Pretraining is conducted for 100,000 steps with a batch size of 8, using a maximum learning rate (LR) of 0.0001, a 10% warm-up phase, and a linear LR decay [12]. The multi-modal transformer adopts 12 layers and a 768-dimensional embedding size, consistent with LayoutLMv3BASE. Pretraining runs for 375,000 steps with a batch size of 4, using a maximum LR of 0.0001, a 4.8% warm-up phase, and linear decay [8]. During finetuning, we use a batch size of 2 and an initial LR of 0.0005, with a performance-based LR scheduler. If validation performance does not improve for 5 epochs, the LR is reduced by 10%; training is terminated after 9 consecutive non-improving epochs. All the models are trained with one NVIDIA A100 (40GB) GPU. [^1] Code and data: https://github.com/knowledge-computing/multimodal-text-linking
@@ 4.4.5 Visualizations
Figure 4: Successful examples with link recall $\geq$90%
@@ Appendix 0.A Complexity Analysis
Method #Param (×$\times 10^{6}$) Time (s/image)
LIGHT - Polygon Encoder $70.1$ $0.16$
BERTBASE $111.0$ $0.17$
LayoutLMv3BASE $130.4$ $0.19$
LIGHT $197.1$ $0.34$
@@ Appendix 0.B Dataset Statistics
Words $34.518$ $5.544$ $128.457$
Valid Words $30.563$ $4.860$ $111.821$
Label Groups $21.205$ $3.502$ $78.582$
Groups with $>$1 Words $9.772$ $1.453$ $36.655$